out of 800 progeny of a three-point testcross there were 24 double crossover recombinants, whereas 30 had been expected on the basis of no interference, the Interference is ____? *The answer is 0.20, (Please show calculations) and how to solve the problem step by step.
Answer #1

Interference = 1 - coefficient of coincidence

Coefficient of coincidence is the ratio of observed to expected double recombinants.

In the above question, he observed number of DCO is 24 and expected number of DCO is 30. Therefore, coincidence of coefficient is 24 / 30 = 0.8.

Since interference is 1 - C. O.C., = 1 - 0.8 = 0.2

Interference for the given three point testcross is 0.2